OpenStudy (anonymous):

What is the value of x? 6x-2(x+4)=12

OpenStudy (anonymous):

Answer: x = 5. I am in a rush so yeah

OpenStudy (anonymous):

Others will help explain why but I have to go

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (denonakavro):

6x-2(x+4) = 12 open parenthesis by multiplying 6x-2x+8 = 12 subtract similar valıes 4x+8 = 12 subtract 8 from both sides 4x = 4 divide both sides by 4 x = 1

OpenStudy (anonymous):

6x-2(X+4)=12 Simplify your parenthesis 6x [-2x + (-8)]=12 Combine you like terms 4x + (-8) = 12 Add 8 to both sides to cancel it out 4x = 20 Divide both sides by 4 to find how much x equals 1x = 5

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

You're welcome. Anymore questions?

OpenStudy (anonymous):

yes should i close this question?

OpenStudy (anonymous):

Yes you should close this questions. (The Moderators would appreciate it) Please link me in your other questions. I am working on my K12 and I may not realize when you post a question.

OpenStudy (anonymous):

ok